# Deep Learning for Camera Calibration and Beyond: A Survey

### v2

@@ 5.4.1 Technique Summary
The above works are devoted to exploring different homography parameterizations such as 4-pt parameterization[28], perspective field[85], and motion bases representation[54], which contributes to better convergence and performance. Other works tend to design various network architectures. In particular, cascaded and iterative solutions are proposed to refine the performance progressively, which can be further combined together to reach higher accuracy. To make the methods more practical, various challenging problems are preliminarily addressed, such as cross resolutions[147], multiple modalities[146, 165], dynamic objects[113, 165], and non-planar scenes[53, 55, 44], etc.
@@ 5.4.2 Challenge and Future Effort
(1) To overcome the first challenge, we can design various strategies to enhance resolution robustness, such as resolution-related data augmentation, and continual learning on multiple datasets with different resolutions. Besides, we can also formulate a resolution-free parameterization form. The perspective field [85] is a typical case, which represents the homography as dense correspondences with the same resolution as input images. But it requires RANSAC as the post-processing approach, introducing extra computational complexity, especially in the case of extensive correspondences. Therefore, a resolution-free and efficient parameterization form should be explored.
@@
(3) As there is no interaction between different image features in the feature extractor, it is reasonable to assume that outlier rejection should occur after feature extraction. It is not possible to identify outliers within a single image as the depth alone cannot be used as an outlier cue. For example, images captured by purely rotated cameras do not contain parallax outliers. Additionally, it seems intuitive to learn the capability of outlier rejection by combining global and local correlation, similar to the insight of RANSAC.
@@ 6 Cross-Sensor Model
In this part, we mainly review the literature on learning-based camera-LiDAR calibration, i.e., predicting the 6-DoF rigid body transformation between a camera and a 3D LiDAR, without requiring any presence of specific features or landmarks in the implementation. Like calibration works on other types of cameras/systems, this research field can also be classified into regression-based solutions and flow/reconstruction-based solutions. But we are prone to follow the special matching principle in camera-LiDAR calibration and divide the existing learning-based literature into three categories: pixel-level solution, semantics-level solution, and object/keypoint-level solution.
@@ 6.1 Pixel-level Solution
The first deep learning technique in camera-LiDAR calibration, RegNet [30], used CNNs to combine feature extraction, feature matching, and global regression to infer the 6-DoF extrinsic parameters. It processed the RGB and LiDAR depth map separately and branched two parallel data network streams. Then, a specific correlation layer was proposed to convolve the stacked LiDAR and RGB features as a joint representation. After this feature matching, the global information fusion and parameter regression were achieved by two fully connected layers with a Euclidean loss function. Motivated by this work, the subsequent works made a further step into more accurate camera-LiDAR calibration in terms of the geometric constraint [86, 131], temporal correlation [131], loss design [130], feature extraction [182], feature matching [155, 135], feature fusion [182], and calibration representation [156, 179].
@@
For example, as shown in Figure 10, CalibNet [86] designed a network to predict calibration parameters that maximize the geometric and photometric consistency of images and point clouds, solving the underlying physical problem by 3D Spatial Transformers [216]. To refine the calibration model, CalibRCNN [131] presented a synthetic view and an epipolar geometry constraint to measure the photometric and geometric inaccuracies between consecutive frames, of which the temporal information learned by the LSTM network has been investigated in the learning-based camera-LiDAR calibration for the first time. Since the output space of the LiDAR-camera calibration is on the 3D Special Euclidean Group () rather than the normal Euclidean space, RGGNet [130] considered Riemannian geometry constraints in the loss function, namely, used a geodesic distance equipped with left-invariant Riemannian metrics to optimize the calibration network. LCCNet [155] exploited the cost volume layer to learn the correlation between the image and the depth transformed by the point cloud. Because the depth map ignores the 3D geometric structure of the point cloud, FusionNet [182] leveraged PointNet++ [217] to directly learn the features from the 3D point cloud. Subsequently, a feature fusion with Ball Query [217] and attention strategy was proposed to effectively fuse the features of images and point clouds.
CFNet [156] first proposed the calibration flow for camera-LiDAR calibration, which represents the deviation between the positions of initial projected 2D points and ground truth. Compared to directly predicting extrinsic parameters, learning the calibration flow helped the network to understand the underlying geometric constraint. To build precise 2D-3D correspondences, CFNet [156] corrected the originally projected points using the estimated calibration flow. Then the efficient Perspective-n-Point (EPnP) algorithm was applied to calculate the final extrinsic parameters by RANSAC. Because RANSAC is nondifferentiable, DXQ-Net [179] further presented a probabilistic model for LiDAR-camera calibration flow, which estimates the uncertainty to measure the quality of LiDAR-camera data association. Then, the differentiable pose estimation module was designed for solving extrinsic parameters, back-propagating the extrinsic error to the flow prediction network.
@@ 6.2 Semantics-level Solution
Semantic features can be well learned and represented by deep neural networks. A perfect calibration enables to accurately align the same instance in different sensors. To this end, some works [134, 132, 160, 45] explored to guide the camera-LiDAR calibration with the semantic information. SOIC [134] calibrated and transforms the initialization issue into the semantic centroids’ PnP problem using semantic information. Since the 3D semantic centroids of the point cloud and the 2D semantic centroids of the picture cannot match precisely, a matching constraint cost function based on the semantic components was presented. SSI-Calib [132] reformulated the calibration as an optimization problem with a novel calibration quality metric based on semantic features. Then, a non-monotonic subgradient ascent algorithm was proposed to calculate the calibration parameters. Other works utilized the off-the-shelf segmentation networks for point cloud and image, and optimized the calibration parameters by minimizing semantic alignment loss in single-direction [160] and bi-direction [45].
@@ 6.3 Object/Keypoint-level Solution
ATOP [181] designed an attention-based object-level matching network, i.e., Cross-Modal Matching Network to explore the overlapped FoV between camera and LiDAR, which facilitated generating the 2D-3D object-level correspondences. 2D and 3D object proposals were detected by YOLOv4 [218] and PointPillar [219]. Then, two cascaded PSO-based algorithms [220] were devised to estimate the calibration extrinsic parameters in the optimization stage. Using the deep declarative network (DDN) [221], RGKCNet [183] combined the standard neural layer and a PnP solver in the same network, formulating the 2D–3D data association and pose estimation as a bilevel optimization problem. Therefore, both the feature extraction capability of the convolutional layer and the conventional geometric solver can be employed. Microsoft’s human keypoint extraction network [222] was applied to detect the 2D–3D matching keypoints. Additionally, RGKCNet [183] presented a learnable weight layer that determines the keypoints involved in the solver, enabling the whole pipeline to be trained end-to-end.
@@ 6.4.1 Technique Summary
The current method can be briefly classified based on the principle of building 2D and 3D matching, namely, the calibration target. In summary, most pixel-level solutions utilized the end-to-end framework to address this task. While these solutions delivered satisfactory performances on specific datasets, their generalization abilities are limited. Semantics-level and object/keypoint-level methods derived from traditional calibration offered both acceptable performances and generalization abilities. However, they heavily relied on the quality of fore-end feature extraction.
@@ 6.4.2 Research Trend
(1) Network architecture is becoming more complex with the use of different structures for feature extraction, matching, and fusion. Current methods employ strategies like multi-scale feature extraction, cross-modal interaction, cost-volume establishment, and confidence-guided fusion.
@@ 7 Benchmark
As there is no public and unified benchmark in learning-based camera calibration, we contribute a dataset that can serve as a platform for generalization evaluations. In this dataset, the images and videos are captured by different cameras under diverse scenes, including simulation environments and real-world settings. Additionally, we provide the calibration ground truth, parameter label, and visual clues in this dataset based on different conditions. Figure 11 shows some samples of our collected dataset. Please refer to the evaluation of representative calibration methods on this benchmark in supplementary material.
@@
Distortion Model. We created a comprehensive dataset for the distortion camera model, with a focus on wide-angle cameras. The dataset is comprised of three subcategories. The first is a synthetic dataset, which was generated using the widely-used 4th order polynomial model. It contains both circular and rectangular structures, with 1,000 distortion-rectification image pairs. The second subcategory consists of data captured under real-world settings, derived from the raw calibration data for around 40 types of wide-angle cameras. For each calibration data, the intrinsics, extrinsics, and distortion coefficients are available. Finally, we exploit a car equipped with different cameras to capture video sequences. The scenes cover both indoor and outdoor environments, including daytime and nighttime footage.
@@ 8.1 Sequences
Bundle adjustment is a well-established technique central to Multi-View Stereo (MVS) and Simultaneous Localization and Mapping (SLAM) using multi-view constraints. Traditional bundle adjustment focuses on pose estimation, often under the assumption of pre-calibrated cameras, thus sidelining the nuances of camera parameter fine-tuning. While learning-based camera calibration has made significant strides, most methods are tailored for a single image. We have highlighted intrinsics calibration to underscore how sequence constraints bolster prediction accuracy. Notably, there is a burgeoning interest in integrating bundle adjustment into end-to-end deep learning pipelines. By transitioning from conventional keypoint extraction and matching to learning-based methods, recent works [228, 229, 230, 231, 232] propose differentiable bundle adjustment layers to refine pose, depth, and camera parameters together. Consequently, there is immense potential in further harnessing sequence constraints for accurate calibration. Current methods combine front-end matching with a back-end solver, which can be inefficient and unreliable in cases like fast motion. We suggest separating front-end and back-end refinements, using large models for features and introducing more trainable parameters in optimization.
@@ 8.3 Pre-training
Pre-training on ImageNet [69] has become a widely used strategy in deep learning. However, recent studies [96] have shown that this approach provides less benefit for specific camera calibration tasks, such as wide-angle camera calibration. This is due to two main reasons: the data gap and the task gap. The ImageNet dataset only contains perspective images without distortions, making the initialized weights of neural networks irrelevant to distortion models. Furthermore, He et al. [233] demonstrated that the task of ImageNet pre-training has limited benefits when the final task is more sensitive to localization. As a result, the performance of extrinsics estimation may be impacted by this task gap. Moreover, pre-training beyond a single image and a single modality, to our knowledge, has not been thoroughly investigated in the related field. We suggest that designing a customized pre-training strategy for learning-based camera calibration is an interesting area of research.
@@ 8.4 Implicit Unified Model
Deep learning-based camera calibration methods use traditional parametric camera models, which lack the flexibility to fit complex situations. Non-parametric camera models relate each pixel to its corresponding 3D observation ray, overcoming parametric model limitations. However, they require strict calibration targets and are more complex for undistortion, projection, and unprojection. Deep learning methods show potential for calibration tasks, making non-parametric models worth revisiting and potentially replacing parametric models in the future. Moreover, they allow for implicit and unified calibration, fitting all camera types through pixel-level regression and avoiding explicit feature extraction and geometry solving. Researchers combined the advantages of implicit and unified representation with the Neural Radiance Field (NeRF) for reconstructing 3D structures and synthesizing novel views. Self-calibration NeRF [191] has been proposed for generic cameras with arbitrary non-linear distortions, and end-to-end pipelines have been explored to learn depth and ego-motion without calibration targets. We believe the implicit and unified camera models could be used to optimize learning-based algorithms or integrated into downstream 3D vision tasks.
@@ 9 Conclusion
In this paper, we present a comprehensive survey of the recent efforts in the area of deep learning-based camera calibration. Our survey covers conventional camera models, classified learning paradigms and learning strategies, detailed reviews of the state-of-the-art approach, a public benchmark, and future research directions. To exhibit the development process and link the connections between existing works, we provide a fine-grained taxonomy that categorizes literature by jointly considering camera models and applications. Moreover, the relationships, strengths, distinctions, and limitations are thoroughly discussed in each category. An open-source repository will keep updating regularly with new works and datasets. We hope that this survey could promote future research in this field.
@@ Acknowledgment
We thank Leidong Qin and Shangrong Yang at Beijing Jiaotong University for the partial dataset collection. We thank Jinlong Fan at the University of Sydney for the insightful discussion.

### v3

@@ 5.4.1 Technique Summary
The above works are devoted to exploring different homography parameterizations such as 4-pt parameterization[29], perspective field[236], and motion bases representation[60], which contributes to better convergence and performance. Others tend to design various network architectures. In particular, cascaded and iterative solutions are proposed to refine the performance progressively, which can be combined jointly to reach higher accuracy. To make the methods more practical, various challenging problems are preliminarily addressed, such as cross resolutions[249], multiple modalities[248, 255, 298, 300], dynamic objects[240, 255], and non-planar scenes[59, 61, 49, 287], etc.
@@ 5.4.2 Challenge and Future Effort
(1) To overcome the first challenge, we can design various strategies to enhance resolution robustness, such as resolution-related data augmentation, and continual learning on multiple datasets with different resolutions. Besides, we can also formulate a resolution-free parameterization form. The perspective field [236] is a typical case, which represents the homography as dense correspondences with the same resolution as input images. But it requires RANSAC as the post-processing approach, introducing extra computational complexity, especially in extensive correspondences. Therefore, a resolution-free and efficient parameterization form could be explored.
@@
(3) As there is no interaction between different image features in the feature extractor, it is reasonable to assume that outlier rejection should occur after feature extraction. Identifying outliers within a single image is impossible as the depth alone cannot be used as an outlier cue. For example, images captured by purely rotated cameras do not contain parallax outliers. Additionally, it seems intuitive to learn the capability of outlier rejection by combining global and local correlation, similar to the insight of RANSAC.
@@ 6 Cross-Sensor Model
In this part, we mainly review the literature on learning-based camera-LiDAR calibration, i.e., predicting the 6-DoF rigid body transformation between a camera and a 3D LiDAR, without requiring any presence of specific features or landmarks in the implementation. Like the works on other types of cameras/systems, this research field can also be classified into regression-based solutions and flow/reconstruction-based solutions. But we are prone to follow the special matching principle in camera-LiDAR calibration and divide the existing learning-based literature into three categories: pixel-level, semantics-level, and object/keypoint-level solutions.
@@ 6.1 Pixel-level Solution
The first deep learning technique in camera-LiDAR calibration, RegNet [31], used CNNs to combine feature extraction, feature matching, and global regression to infer the 6-DoF extrinsic parameters. It processed the RGB and LiDAR depth maps separately and branched two parallel data network streams. Then, a specific correlation layer was proposed to convolve the stacked LiDAR and RGB features as a joint representation. Subsequently, the global information fusion and parameter regression were achieved by two fully connected layers with an Euclidean loss function. Motivated by this work, the following works made a further step into more accurate calibration in terms of the geometric constraint [237, 243], temporal correlation [243], loss design [242], feature extraction [258], feature matching [252, 247], feature fusion [258], and calibration representation [253, 256], etc.
@@
For example, in Figure 10, CalibNet [237] designed a network to predict calibration parameters that maximize the geometric and photometric consistency of images and point clouds, solving the underlying physical problem by 3D Spatial Transformers [376]. To refine the calibration model, CalibRCNN [243] presented a synthetic view and an epipolar geometry constraint to measure the photometric and geometric inaccuracies between consecutive frames, of which the temporal information learned by the LSTM network has been investigated in the learning-based camera-LiDAR calibration. Since the output space of the LiDAR-camera calibration is on the 3D Special Euclidean Group () rather than the normal Euclidean space, RGGNet [242] considered Riemannian geometry constraints in the loss function, namely, used an geodesic distance equipped with left-invariant Riemannian metrics to optimize the calibration network. LCCNet [252] exploited the cost volume layer to learn the correlation between the image and the depth transformed by the point cloud. Because the depth map ignores the 3D geometric structure of the point cloud, FusionNet [258] leveraged PointNet++ [377] to directly learn the features from the 3D point cloud. Subsequently, a feature fusion with Ball Query [377] and attention strategy was proposed to effectively fuse the features of images and point clouds. CFNet [253] first proposed the calibration flow for camera-LiDAR calibration, which represents the deviation between the positions of initial projected 2D points and ground truth. Compared to directly predicting extrinsic parameters, learning the calibration flow helps the network understand the underlying geometric constraint. To build precise 2D-3D correspondences, CFNet [253] corrected the originally projected points using the estimated calibration flow. Then the efficient Perspective-n-Point (EPnP) algorithm was applied to calculate the extrinsic parameters by RANSAC. Because RANSAC is nondifferentiable, DXQ-Net [256] presented a probabilistic model to estimate the uncertainty to measure the quality of LiDAR-camera data association. Then, the differentiable pose estimation module was designed for solving extrinsic parameters, back-propagating the extrinsic error to the flow prediction network. Recent efforts further improve the calibration performance by presenting/introducing cross-modal graph neural network [277], multi-head correlation module [316], vanishing point and horizon line estimation [322], monocular depth estimation [270], homogeneous local-global aware representation [319], and hierarchical and iterative feature matching [321], etc.
@@ 6.2 Semantics-level Solution
Semantic features can be well learned and represented by neural networks. A perfect calibration enables to accurately align the same instance in different sensors. To this end, some works [246, 244, 254, 50, 268, 323] explored to guide the calibration with the semantic information. SOIC [246] calibrated and transformed the initialization issue into the semantic centroids’ PnP problem. Since the 3D semantic centroids of the point cloud and the 2D semantic centroids of the image cannot match precisely, a matching constraint cost function based on the semantic components was presented. SSI-Calib [244] reformulated the calibration as an optimization problem with a novel calibration quality metric based on semantic features. Then, a non-monotonic subgradient ascent was proposed to calculate the calibration parameters. Other works used the off-the-shelf segmentation networks [320] and optimized the calibration parameters by minimizing semantic alignment loss in single-direction [254] and bi-direction [50].
@@ 6.3 Object/Keypoint-level Solution
ATOP [257] designed an attention-based object-level matching network, i.e., Cross-Modal Matching Network to explore the overlapped FoV between camera and LiDAR, which facilitated generating the 2D-3D object-level correspondences. 2D and 3D object proposals were detected by YOLOv4 [378] and PointPillar [379]. Then, two cascaded PSO-based algorithms [380] were devised to estimate the calibration extrinsic parameters in the optimization stage. Using the deep declarative network (DDN) [381], RGKCNet [259] combined the standard neural layer and a PnP solver in the same network, formulating the 2D–3D data association and pose estimation as a bilevel optimization problem. Thus, both the feature extraction capability of the convolutional layer and the conventional geometric solver can be employed. Additionally, RGKCNet [259] presented a learnable weight layer that determines the keypoints involved in the solver, enabling the whole pipeline to be trained end-to-end. P2O-Calib [275] proposed a target-less calibration approach using the 2D-3D edge point extraction, which effectively investigates the occlusion relationship in 3D space.
@@ 6.4.1 Technique Summary
The current method can be briefly classified based on the principle of building 2D and 3D matching, namely, the calibration target. In summary, most pixel-level solutions utilized the end-to-end framework to address this task. While these solutions delivered satisfactory performances on specific datasets, their generalization abilities are limited. Semantics-level and object/keypoint-level methods derived from traditional calibration offered both acceptable performances and generalization abilities. However, they heavily relied on the quality of fore-end feature extraction [320]. Recent efforts [272, 312, 314] jointly learn the cross-modal calibration and the implicit scene representation [382] using differentiable volume rendering, showing more scalable applications than existing calibration solutions.
@@ 6.4.2 Research Trend
(1) Network is becoming more complex with the use of different structures for feature extraction, matching, and fusion. Current methods employ strategies like multi-scale feature extraction, cross-modal interaction, cost-volume establishment, and confidence-guided fusion.
@@ 7 Benchmark
As there is no public and unified benchmark in learning-based camera calibration, we contribute a dataset that can serve as a platform for generalization evaluations. In this dataset, the images and videos are captured by different cameras under diverse scenes, including simulation environments and real-world settings. Besides, we provide the calibration ground truth, parameter label, and visual clues based on different conditions. Figure 11 shows some samples of our collected dataset. Please refer to the evaluation of representative calibration methods on this benchmark in supplementary material.
@@
Distortion Model. We created a comprehensive dataset for the distortion camera model, with a focus on wide-angle cameras. The dataset is comprised of three subcategories. The first is a synthetic dataset, which was generated using the widely-used 4th order polynomial model. It contains both circular and rectangular structures, with 1,000 distortion-rectification image pairs. The second consists of data captured under real-world settings, derived from the raw calibration data for around 40 types of wide-angle cameras. For each calibration data, the intrinsics, extrinsics, and distortion coefficients are provided. Finally, we exploited a car equipped with different cameras to capture video sequences. The scenes cover both indoor and outdoor, including daytime and nighttime footage.
@@ 8.1 Sequences
Bundle adjustment is a well-established technique central to Multi-View Stereo (MVS) and Simultaneous Localization and Mapping (SLAM) using multi-view constraints. Traditional bundle adjustment focuses on pose estimation, often under the assumption of pre-calibrated cameras, thus sidelining the nuances of camera parameter fine-tuning. While learning-based camera calibration has made significant strides, most methods are tailored for a single image. We have highlighted intrinsics calibration to underscore how sequence constraints bolster prediction accuracy. Notably, there is a burgeoning interest in integrating bundle adjustment into end-to-end deep learning pipelines. By transitioning from conventional keypoint extraction and matching to learning-based methods, recent works [387, 388, 389, 390, 174, 337, 391] propose differentiable bundle adjustment layers to refine pose, depth, and camera parameters together. Consequently, there is immense potential in further harnessing sequence constraints for accurate calibration. Current methods combine front-end matching with a back-end solver, which can be inefficient and unreliable in cases like fast motion. We suggest separating front-end and back-end refinements, using large models for features and introducing more trainable parameters in optimization.
@@ 8.3 Pre-training
Pre-training on ImageNet [76] has become a widely used strategy in deep learning. However, recent studies [97] have shown that this approach provides less benefit for specific camera calibration tasks, such as wide-angle camera calibration. This is due to two main reasons: the data gap and the task gap. The ImageNet dataset only contains perspective images without distortions, making the initialized weights of networks irrelevant to distortion models. Furthermore, He et al. [392] demonstrated that the task of ImageNet pre-training has limited benefits when the final task is more sensitive to localization. As a result, the performance of extrinsics estimation may be impacted by this task gap. Moreover, pre-training beyond a single image and a single modality, to our knowledge, has not been investigated in the related field. We suggest that designing a customized pre-training strategy for camera calibration is an interesting area of research.
@@ 8.4 Implicit Unified Model
Deep learning-based camera calibration methods use traditional parametric camera models, which lack the flexibility to fit complex situations. Non-parametric camera models relate each pixel to its corresponding 3D observation ray, overcoming parametric model limitations. However, they require strict calibration targets and are more complex for undistortion, projection, and unprojection. Deep learning methods show potential for calibration tasks, making non-parametric models worth revisiting and potentially replacing parametric models. Moreover, they allow for implicit and unified calibration, fitting all camera types through pixel-level regression and avoiding explicit feature extraction and geometry solving. Researchers combined the advantages of implicit and unified representation with the Neural Radiance Field (NeRF) for reconstructing 3D structures and synthesizing novel views. Self-calibration NeRF [352] has been proposed for generic cameras with arbitrary non-linear distortions, and end-to-end pipelines have been explored to learn depth and ego-motion without calibration targets. We believe the implicit and unified camera models could be used to optimize learning-based algorithms or integrated into downstream 3D vision tasks.
@@ 9 Conclusion
In this paper, we present a comprehensive survey of the recent efforts in deep learning-based camera calibration. Our survey covers conventional camera models, classified learning paradigms and learning strategies, detailed reviews of the state-of-the-art approach, a public benchmark, and future research directions. To exhibit the development process and link the connections between existing works, we provide a fine-grained taxonomy that categorizes literature by jointly considering camera models and applications. Moreover, the relationships, strengths, distinctions, and limitations are thoroughly discussed in each category. An open-source repository will keep updating regularly with new works and datasets. We hope that this survey could promote future research in this field.
@@ Acknowledgment
We thank Leidong Qin and Shangrong Yang at Beijing Jiaotong University for the partial dataset collection. We thank Jinlong Fan at the University of Sydney for the insightful discussion. We appreciate all reviewers for their insightful comments and invaluable suggestions.